Find the exact value of each trig expression
Tan 135 - tan 15
----------------------
1 + tan 135 tan 15

Answer by stanbon(75887)

Find the exact value of each trig expression
Tan 135 - tan 15
----------------------
1 + tan 135 tan 15
==========================
= tan(135-15)
-----
= tan(120) = -tan(30) = -1/sqrt(3)
================
Cheers,
Stan H.
----------------